# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v1

@@ 3.1 Mathematical Formalism
$\hat{\theta}=\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{n=1}^{N}\ell(\mathcal{G}_{n},\theta),\quad\hat{\theta}_{aug}=\operatorname*{arg\,min}_{\theta}\frac{1}{N\times M}\sum_{n=1}^{N}\sum_{m=1}^{M}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta).$
@@ 3.4 Analyzing the Generalization Ability of the Augmented Graphs via Influence Functions
Given a test graph $\mathcal{G}_{k}$ from the test set, let $\hat{\theta}=\arg\min_{\theta}\mathcal{L}$ be the GNN parameters that minimize the objective function in Eq. 3. The impact of upweighting the objective function $\mathcal{L}$ to $\mathcal{L}_{i,j}^{\text{aug}}=\mathcal{L}+\epsilon_{n,m}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)$, where $\mathcal{G}_{n}^{\lambda_{n,m}}$ is an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter, on the model performance on the test graph $\mathcal{G}_{k}^{test}$ is given by
@@ 5 Conclusion
Appendix: Gaussian Mixture Models Based Augmentation Enhances GNN Generalization
@@
For the term $u_{4}$, we apply McDiarmid’s inequality. Since the classification loss satisfy $\ell(\cdot)\in[0,1]$, we get,
$\forall\{(\mathcal{G}_{n},y_{n})\}_{n=1}^{N},\{(\mathcal{G^{\prime}}_{n},y^{\prime}_{n})\}_{n=1}^{N},\theta,$ such that $\forall k\neq n,\quad\mathcal{G}_{n}=\mathcal{G^{\prime}}_{n}$ and $\forall k=n,\quad\mathcal{G}_{k}\neq\mathcal{G^{\prime}}_{k}$:
$\left|\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n},\theta)\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$ $=\left|\frac{1}{N}\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\left[\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)\right]-\frac{1}{N}\sum_{n=1}^{N}\frac{1}{M}\sum_{m=1}^{M}\left[\ell(\mathcal{G^{\prime}}_{n}^{\lambda_{n,m}},\theta)\right]\right|$
$=\left|\frac{1}{N\times M}\sum_{m=1}^{M}\sum_{n=1}^{N}\left[\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)-\ell(\mathcal{G^{\prime}}_{n}^{\lambda_{n,m}},\theta)\right]\right|$
$=\left|\frac{1}{N\times M}\sum_{m=1}^{M}\left[\ell(\mathcal{G}_{k}^{\lambda_{k,m}},\theta)-\ell(\mathcal{G^{\prime}}_{k}^{\lambda_{k,m}},\theta)\right]\right|$
$\leq\frac{1}{N\times M}\times 2\times M$
@@ Appendix D Ablation Study
In Table 5, we compared the performance of two approaches to fit the GMM on the graph representations of each class. We particularly compared the EM algorithm, discussed in the main paper, and the Variational Bayesian Estimation (VB) technique. Bishop & Nasrabadi (2006). As noticed in the results, the two methods lead to approximatly the same results.
Table 5: Ablation Study, VB: Variational Bayesian estimation.
@@
GMMv2 w/ EM 71.00 (4.40) 49.82 (4.26) 76.05 (6.47) 70.97 (5.07) 71.90 (2.81)
GMMv2 w/ VB 71.00 (4.21) 49.53 (4.26) 76.05 (6.47) 70.97 (4.52) 71.64 (2.90)

### v2

@@ 3.1 Mathematical Formalism
$\hat{\theta}$ $=\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{n=1}^{N}\ell(\mathcal{G}_{n},\theta),$
$\hat{\theta}_{aug}$ $=\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta)\right]\simeq\operatorname*{arg\,min}_{\theta}\frac{1}{N\times M}\sum_{n=1}^{N}\sum_{m=1}^{M}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta).$
@@
The findings of Theorem 3.1 hold for all norms defined on the graph input space. Specifically, let us consider the graph space $(\mathbb{G},\lVert\cdot\rVert_{\mathbb{G}})$ and the feature space $(\mathbb{X},\lVert\cdot\rVert_{\mathbb{X}})$, where $\lVert\cdot\rVert_{\mathbb{G}}$ and $\lVert\cdot\rVert_{\mathbb{X}}$ denote the norms applied to the graph structure and features, respectively. Assuming a maximum number of nodes per graph, which is a realistic assumption for real-world data, the product space $\mathbb{G}\times\mathbb{X}$ is a finite-dimensional real vector space, and all the norms are equivalent. Thus, the choice of norm does not affect the theorem, as long as the Lipschitz constant is adjusted accordingly. Additional details and insights on the graph distance metrics can be found in Appendix G.
@@ 3.4 Analyzing the Generalization Ability of the Augmented Graphs via Influence Functions
Given a test graph $\mathcal{G}_{k}$ from the test set, let $\hat{\theta}=\arg\min_{\theta}\mathcal{L}$ be the GNN parameters that minimize the objective function in Eq. 3. The impact of upweighting the objective function $\mathcal{L}$ to $\mathcal{L}_{n,m}^{\text{aug}}=\mathcal{L}+\epsilon_{n,m}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)$, where $\mathcal{G}_{n}^{\lambda_{n,m}}$ is an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter, on the model performance on the test graph $\mathcal{G}_{k}^{test}$ is given by
@@ 5 Conclusion
For the term $u_{4}$, we apply McDiarmid’s inequality. Since the classification loss satisfy $\ell(\cdot)\in[0,1]$, we get for $k\in\{0,\ldots,N\}$,
$\forall\{(\mathcal{G}_{n},y_{n})\}_{n=1}^{N},\{(\mathcal{G^{\prime}}_{n},y^{\prime}_{n})\}_{n=1}^{N},\theta,$ such that $\forall n\neq k,\quad\mathcal{G}_{n}=\mathcal{G^{\prime}}_{n}$ and $\mathcal{G}_{k}\neq\mathcal{G^{\prime}}_{k}$:
$\left|\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n},\theta)\right]-\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$ $=\frac{1}{N}\left|\sum_{n=1}^{N}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n}^{\prime},\theta)\right]\right|$
$\leq\frac{1}{N}\left|\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{k},\theta)-\ell(\mathcal{G}_{k}^{\prime},\theta)\right]\right|$
@@
The first equality is obtained by your claim that $\forall n\neq k,\quad\mathcal{G}_{n}=\mathcal{G^{\prime}}_{n}$ and $\mathcal{G}_{k}\neq\mathcal{G^{\prime}}_{k}$, the last inequality is obtained by the fact that $\ell(\cdot)\in[0,1]$.
@@ Appendix D Ablation Study
To provide additional comparison and motivate the use of GMMs with the EM algorithm, we expanded our evaluation to include additional methods for modeling the distribution of the graph representations. Specifically, the comparison includes:
  • GMM w/ Variational Bayesian Inference (VBI): We specifically compared the Expectation-Maximization (EM) algorithm, discussed in the main paper, with the Variational Bayesian (VB) estimation technique for parameter estimation of each Gaussian Mixture Model (GMM) (Tzikas et al., 2008) for both the GCN and GIN models. The objective of including this baseline is to explore alternative approaches for fitting GMMs to the graph representations.
  • Kernel Density Estimation (KDE) : KDE is a Neighbor-Based Method and a non-parametric approach to estimating the probability density (Härdle et al., 2004). KDE estimates the probability density function by placing a kernel function (e.g., Gaussian) at each data point. The sum of these kernels approximates the underlying distribution. Sampling can be done using techniques like Metropolis-Hastings. The purpose of using KDE as a baseline is to evaluate alternative distributions different from the Gaussian Mixture Model (GMM).
  • Copula-Based Methods: We model the dependence structure between variables using copulas, while marginal distributions are modeled separately. We sample from marginal distributions and then transform them using the copula (Nelsen, 2006).
  • Generative Adversarial Network (GAN): GANs are powerful generative models that learn to approximate the data distribution through an adversarial process between two neural networks. To evaluate the performance of deep learning-based generative approaches for modeling graph representations, we included tGAN, a GAN architecture specifically designed for tabular data (Yang et al., 2012). We particularly train tGAN on the graph representations and then sample new graph representations from the generator.
Table 5: Ablation Study on the density estimation scheme for learned GCN representations.
@@
GMM w/ EM 71.00 (4.40) 49.82 (4.26) 76.05 (6.47) 70.97 (5.07) 71.90 (2.81)
GMM w/ VBI 71.00 (4.21) 49.53 (4.26) 76.05 (6.47) 70.97 (4.52) 71.64 (2.90)
KDE 55.90 (10.29) 39.53 (2.87) 66.64 (6.79) 59.56 (2.62) 58.66 (3.97)
Copula 69.80 (4.04) 47.13 (3.45) 74.44 (6.26) 65.04 (3.37) 65.70 (3.04)
GAN 70.60 (3.41) 48.80 (5.51) 75.52 (4.96) 69.98 (5.46) 66.26 (3.72)
Table 6: Ablation Study on the density estimation scheme for learned GIN representations.
Model IMDB-BINARY IMDB-MULTI MUTAG PROTEINS DD
GMM w/ EM 71.70 (4.24) 49.20 (2.06) 88.83 (5.02) 71.33 (5.04) 68.61 (4.62)
GMM w/ VBI 71.40 (2.65) 47.80 (2.22) 88.30 (5.19) 70.25 (4.65) 67.82 (4.96)
KDE 69.10 (3.93) 41.46 (3.02) 77.60 (6.83) 60.37 (3.04) 67.48 (6.18)
Copula 70.60 (2.61) 47.60 (2.29) 88.30 (5.19) 70.16 (4.55) 67.91 (4.90)
GAN 70.50 (3.80) 48.40 (1.71) 88.83 (5.02) 71.33 (5.55) 67.74 (4.82)
We compare these approaches for both the GCN and GIN models in Tables 5 and 6, respectively. As noticed, GMM with EM consistently outperforms the alternative methods across most datasets in terms of accuracy. The VBI method, an alternative approach for estimating GMM parameters, yields comparable performance to the EM algorithm. This consistency across datasets highlights the effectiveness and robustness of GMMs in capturing the underlying data distribution.
In certain cases, particularly with the GIN model, we observed competitive performance from the GAN approach, which, unlike GMM, requires additional training. Hence, GMMs provide a more straightforward and efficient solution.
@@ Appendix E Training and Augmentation time
We compare the data augmentation times of our approach and the baselines in Table 7. In addition to outperforming the baselines on most datasets, our approach offers an advantage in terms of time complexity. The training time of baseline models varies depending on the augmentation strategy used, specifically, whether it involves pairs or individual graphs. Even in cases where a graph augmentation has a low computational cost for some baselines, training can still be time-consuming as multiple augmented graphs are required to achieve satisfactory test accuracy. For instance, methods like DropEdge, DropNode, and SubMix, while computationally simple, require generating multiple augmented samples at each epoch, thereby increasing the overall training time. In contrast, GMM-GDA introduces a more efficient approach by generating only one augmented graph per training instance, which is reused across all epochs. This design ensures a balance between computational efficiency and augmentation effectiveness, reducing the overall training burden while maintaining strong performance. The only baseline that is more time-efficient than our approach is GeoMix; however, our method consistently outperforms GeoMix across all settings, as shown in Tables 1 and 2.
@@ F.2 Implementation Details
Appendix G Graph Distance Metrics
Let us consider the graph space $(\mathbb{G},\lVert\cdot\rVert_{\mathbb{G}})$ and the feature space $(\mathbb{X},\lVert\cdot\rVert_{\mathbb{X}})$, where $\lVert\cdot\rVert_{\mathbb{G}}$ and $\lVert\cdot\rVert_{\mathbb{X}}$ denote the norms applied to the graph structure and features, respectively. When considering only structural changes, with fixed node features, the distance between two graphs $\mathcal{G}^{\lambda},\mathcal{G}$ is defined as
$\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|=\lVert A-A^{\lambda}\rVert_{\mathbb{G}},$ (8)
where $A^{\lambda},A$ are respectively the adjacency matrix of $\mathcal{G}^{\lambda},\mathcal{G},$ and the norm $\lVert\cdot\rVert_{\mathbb{G}}$ can be for example the Frobenius or spectral norm. If both structural and feature changes are considered, the distance extends to:
$\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|=\alpha\lVert A-A^{\lambda}\rVert_{\mathbb{G}}+\beta\lVert X-X^{\lambda}\rVert_{\mathbb{X}},$ (9)
where $X^{\lambda},X$ are the node feature matrices of $\mathcal{G}^{\lambda},\mathcal{G}$ respectively, and $\alpha,\beta$ are hyperparameters controlling the contribution of structural and feature differences.
In most baselines graph augmentation techniques, such as for instance $\mathcal{G}$-Mixup, SubMix, and DropNode, the alignment between nodes in the original graph $\mathcal{G}$ and the augmented graph $\mathcal{G}^{\lambda}$ is known. However, in cases where the node alignment is unknown, we must take into account node permutations. The distance between the two graphs is then defined as
$\left\|\mathcal{G}^{\lambda}-\mathcal{G}\right\|=\min_{P\in\Pi}\left(\alpha\lVert A-PA^{\lambda}P^{T}\rVert_{\mathbb{G}}+\beta\lVert X-PX^{\lambda}\rVert_{\mathbb{X}}\right),\qquad\qquad$ (10)
where $\Pi$ is the set of permutation matrices. The matrix $P$ corresponds to a permutation matrix used to order nodes from different graphs. By using Optimal Transport, we find the minimum distance over the set of permutation matrices, which corresponds to the optimal matching between nodes in the two graphs. This formulation represents the general case of graph distance, which has been used in the literature (Abbahaddou et al., 2024).